OpenStudy (anonymous):

Which functions have a positive rate of change and which have a negative rate of change? Column A Column B A. Positive B. Negative 1. 2x − 5y = −15 2. y = −4x + 9 3. x y 2 –10 4 –16 6 –22 4. x y –3 15 –2 11 –1 7

OpenStudy (sleepyjess):

Everything is good on that one except #3 \(\dfrac{-16+10}{4-2}\\ -\dfrac{6}2\)

OpenStudy (anonymous):

ohh I see ty

OpenStudy (sleepyjess):

No problem :)

OpenStudy (sleepyjess):

B is correct :)

OpenStudy (anonymous):

Okay, I just wanna make sure my answers are correct. I got A Which function has the greatest starting value? A. x y 1 −3 2 −8 3 −13 B. y = 3x – 8 C. 8x + 4y = 16 D. A linear function that goes through point 2, 0 and point 0, 2

OpenStudy (sleepyjess):

Hmm.... I've never heard of a starting value for a function.... @jigglypuff314 , can you help?

jigglypuff314 (jigglypuff314):

sorry for the late reply (I was afk) starting value should refer to the y-intercept so which option has the biggest y-intercept? :P

OpenStudy (anonymous):

I got the answer lol , ty for your help